1. A block of mass 2 kg is placed on rough horizontal surface ( coefficient of friction =0.2) and is pulled by horizontal force F = 2t N where t : time in sec. Velocity of block at t =4 sec is-

Grade:12

1 Answers

dee yy tutu
14 years ago
f=umg=4 N 2t=4 at t=2 block will start moving after t=2 net force=2t-4 a=t-2 put a=dV/dT and integrate from t=2 to t=4